a research article reported difference among various therapies in addressing anxiety disorders. F(4,45) = 7.53< .01
assuming equal sample sizes for each group, how many patients participated in each therapy
Answer #1

Solution:

Given:

F(4,45)=7.53

Equal sample sizes in each group.

We have to find number of observations in each group.

That is n=......?

k = number of groups

N= total observations

We have:

Df​Between = k-1

4=k-1

k=4+1

k=5

And

Df​​​​​Within =N-k

45 =N-5

N=45+5

N=50

Thus

n = N / k

n = 50 / 5

n = 10

Thus number of observations in each group = 10